OpenStudy (anonymous):

An acute angle θ is in a right triangle with cos θ = five sixths. What is the value of sec θ?

OpenStudy (campbell_st):

sec = 1/cos fo just find the reciprocal of cos

OpenStudy (anonymous):

square root of eleven divided by six five divided by square root of eleven six fifths eleven divided by five

OpenStudy (anonymous):

which one?

OpenStudy (campbell_st):

well look at it this way  \[\sec ( \theta) = \frac{1}{\cos(\theta)}\] are you sure you have the correct information, the ratio is cos and you need to find sec..?

OpenStudy (anonymous):

Yes I am sure that its one of them

OpenStudy (campbell_st):

because if you know cos \[\cos(\theta) = \frac{5}{6} \] then \[\sec(\theta) = \frac{1}{\cos(\theta)} = \frac{1}{\frac{5}{6}}\] just simplify

OpenStudy (anonymous):

So the answer is 5/6?

OpenStudy (campbell_st):

nope the rule for dividing by a fraction is take the reciprocal and multiply

OpenStudy (anonymous):

so its 6/5 right?

OpenStudy (campbell_st):

so you get \[\frac{1}{\frac{5}{6}} = 1 \times \frac{6}{5} \]

OpenStudy (campbell_st):

correct

OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ = seven eighths. What is the value of cot θ?

OpenStudy (usukidoll):

draw a triangle and then use the a^2+b^2 = c^2 to find the missing value.

OpenStudy (usukidoll):